# Efficient Error and Variance Estimation

### v3

@@ 4.2 Randomized SVD
    Draw a test matrix $\boldsymbol{\Omega}\in\mathbb{K}^{d_{2}\times s}$ with iid isotropic columns.
@@ 5.1 Experimental setup
To evaluate our diagnostics for matrices with different spectral characteristics, we consider synthetic test matrices from [27, §5]:
@@ 5.2 Leave-one-out error estimator for randomized SVD
We set $q=0$. Figure 3 shows the results for three examples in the previous section. In all cases, error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ tracks the true error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{A})$ closely. Additional examples and analogous plots for randomized Nyström approximation are provided in Appendix SM2.
@@ 5.3 Matrix jackknife for projectors onto singular subspaces
Consider the task of computing the projector $\boldsymbol{\Pi}$ onto the dominant 55-dimensional right singular subspace to a matrix $\boldsymbol{A}$. The randomized SVD $\boldsymbol{A}\approx\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{*}$ yields the approximation
@@
Figure 4 shows the mean error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{\Pi})$, standard deviation $\operatorname{Std}(\boldsymbol{X})$, and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ for the same three test matrices as previous. We again set $q=0$. Consistent with Theorem 3.1, the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ is an overestimate of $\operatorname{Std}(\boldsymbol{X})$ by a factor of $2\times$ to $8\times$. While the jackknife is not quantitatively sharp, it provides an order of magnitude estimate of the standard deviation and is a useful diagnostic for the quality of the computed output. Additional examples and plots for randomized Nyström approximations of projectors onto invariant subspaces are provided in Appendix SM2.
@@ 6 Extension: Variance estimates for higher Schatten norms
$\boldsymbol{\widehat{\operatorname{Var}}}_{1}(\boldsymbol{X}_{s-1})\coloneqq\frac{1}{2}\sum_{j=1}^{s-1}\left|\boldsymbol{X}^{(s)}-\boldsymbol{X}^{(j)}\right|^{2}\quad\textrm{and}\quad\boldsymbol{\widehat{\operatorname{Var}}}_{2}(\boldsymbol{X}_{s-1})\coloneqq\frac{1}{2}\sum_{j=1}^{s-1}\left|\boldsymbol{X}^{(s)}-\boldsymbol{X}^{(j)}\right|_{*}^{2}.$
Define the Schatten $p$-norm variance estimate
@@
A matrix generalization of the Efron–Stein–Steele inequality [18, Thm. 4.2] shows this jackknife variance estimate overestimates the Schatten $p$-norm variance in the sense
@@ Appendix A Derivation of update formulas
Assume without loss of generality that $j=s$, instate the notation of Section 4.1, and assume $\boldsymbol{H}$ is invertible. The key ingredient is the following consequence of the Banachiewicz inversion formula [19, eq. (0.7.2)]:
@@
In this section, we discuss additional implementation details for fast computation of the jackknife variance estimate. MATLAB R2022b implementations are provided in Appendix SM3.
@@
Section SM2 Additional numerical experiments
@@
  • Kernel. The QM9 kernel matrix from Section $1.$
@@
Results for the leave-one-out error estimator for the randomized SVD and randomized Nyström approximation appear in Figs. SM1 and SM2. Jackknife variance estimates for projectors onto singular subspaces computed by the randomized SVD and projectors onto invariant subspaces computed by randomized Nyström approximation appear in Figs. SM3 and SM4.
@@
Figure SM2: Leave-one-out error estimator for randomized Nyström. Error and error estimate for randomized Nyström approximation ($q=0$).
@@
Figure SM3: Matrix jackknife for projectors onto singular subspaces. Error, standard deviation, and jackknife estimate for randomized SVD ($q=0$) approximation $\boldsymbol{X}$ to the projector $\boldsymbol{\Pi}$ onto the span of the five dominant right singular vectors Eq. 19. Includes more test matrices than Fig. 4.
@@
Section SM3 MATLAB implementations
In this section, we provide MATLAB R2022b implementations of the leave-one-out error estimate and jackknife variance estimate for the Nyström approximation (Section SM3.1). and randomized SVD (Section SM3.2).
SM3.1 Nyström approximation

### v4

@@ 4.2 Randomized SVD
    Draw a test matrix $\boldsymbol{\Omega}\in\mathbb{C}^{d_{2}\times s}$ with iid isotropic columns.
@@ 5.1 Experimental setup
To evaluate our diagnostics for matrices with different spectral characteristics, we consider synthetic test matrices from [28, §5]:
@@ 5.2 Leave-one-out error estimator for randomized SVD
We set $q=0$. Figure 3 shows the results for three examples in the previous section. In all cases, error estimate $\widehat{\operatorname{Err}}(\boldsymbol{X},\boldsymbol{A})$ tracks the true error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{A})$ closely. Additional examples and analogous plots for randomized Nyström approximation are provided in Appendix SM3.
@@ 5.3 Matrix jackknife for projectors onto singular subspaces
Consider the task of computing the projector $\boldsymbol{\Pi}$ onto the dominant $5$-dimensional right singular subspace to a matrix $\boldsymbol{A}$. The randomized SVD $\boldsymbol{A}\approx\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{*}$ yields the approximation
@@
Figure 4 shows the mean error $\operatorname{Err}(\boldsymbol{X},\boldsymbol{\Pi})$, standard deviation $\operatorname{Std}(\boldsymbol{X})$, and the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ for the same three test matrices as previous. We again set $q=0$. Consistent with Theorem 3.1, the jackknife estimate $\operatorname{Jack}(\boldsymbol{X})$ is an overestimate of $\operatorname{Std}(\boldsymbol{X})$ by a factor of $2\times$ to $8\times$. While the jackknife is not quantitatively sharp, it provides an order of magnitude estimate of the standard deviation and is a useful diagnostic for the quality of the computed output. Additional examples and plots for randomized Nyström approximations of projectors onto invariant subspaces are provided in Appendix SM3.
@@ 6 Extension: Variance estimates for higher Schatten norms
$\boldsymbol{\widehat{\operatorname{Var}}}_{1}(\boldsymbol{X}_{s-1})\coloneqq\frac{1}{2}\sum_{j=1}^{s-1}\left|\boldsymbol{X}^{(s)}-\boldsymbol{X}^{(j)}\right|^{2}\quad\textrm{and}\quad\boldsymbol{\widehat{\operatorname{Var}}}_{2}(\boldsymbol{X}_{s-1})\coloneqq\frac{1}{2}\sum_{j=1}^{s-1}\left|\left(\boldsymbol{X}^{(s)}-\boldsymbol{X}^{(j)}\right)^{*}\right|^{2}.$
Here, $|\cdot|$ denotes the matrix modulus $|\boldsymbol{B}|\coloneqq(\boldsymbol{B}^{*}\boldsymbol{B})^{1/2}$. Define the Schatten $p$-norm variance estimate
@@
A matrix generalization of the Efron–Stein–Steele inequality [19, Thm. 4.2] shows this jackknife variance estimate overestimates the Schatten $p$-norm variance in the sense
@@ Appendix A Derivation of update formulas
Assume without loss of generality that $j=s$, instate the notation of Section 4.1, and assume $\boldsymbol{H}$ is invertible. The key ingredient is the following consequence of the Banachiewicz inversion formula [20, eq. (0.7.2)]:
@@
Acknowledgements
We thank Robert Webber for his advice and feedback.
Disclaimer
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
@@ Acknowledgements
In this section, we discuss additional implementation details for fast computation of the jackknife variance estimate. MATLAB R2022b implementations are provided in Appendix SM4.
@@
Section SM2 Failure of the Efron bootstrap variance estimate for the randomized SVD
The matrix jackknife produces an estimate of the standard deviation that is positively biased (Theorem 3.1)—by up to an order of magnitude in our experiments. It is natural to wonder: Could the bootstrap do better? In this section, we demonstrate that a staightforward application of the bootstrap variance estimate [6, §5.1] to the Halko–Martinsson–Tropp randomized SVD can go badly wrong.
To focus the discussion, consider the task of estimating the standard deviation of the largest singular value $S\coloneqq\sigma_{\rm max}(\boldsymbol{X})$ estimated from the randomized SVD approximation $\boldsymbol{X}=\boldsymbol{X}(\boldsymbol{\Omega})$ (with no subspace iteration, $q=0$). A straightforward application of the classical bootstrap variance estimate of Efron [6, §5.1] would proceed as follows: Fix a number of bootstrap iterations $B\geq 1$. For $j=1,\ldots,B$, do the following:
  1. 1.
    Sample $\boldsymbol{\omega}_{1\star},\ldots,\boldsymbol{\omega}_{s\star}$ uniformly and with replacement from $\{\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{s}\}$.
  2. 2.
[…]
Finally, we define the Efron bootstrap variance estimate as
$\mathrm{Boot}^{2}(S)\coloneqq\frac{1}{B-1}\sum_{j=1}^{B}\left(S^{(j)}-S^{(\cdot)}\right)^{2}\quad\text{where}\quad S^{(\cdot)}=\frac{1}{B}\sum_{j=1}^{B}S^{(j)}.$
[…]
To demonstrate this, consider the diagonal matrix
$\boldsymbol{A}=\operatorname{diag}\left(1,0.99,0.98,\ldots,0.26,\frac{0.25}{1^{2}},\frac{0.25}{2^{2}},\frac{0.25}{3^{2}},\ldots,\frac{0.25}{925^{2}}\right)\in\real^{1000\times 1000}.$
The first 75 singular values of $\boldsymbol{A}$ decay quite slowly, after which they decay at a polynomial rate. We consider the estimate of the largest singular value $S=\sigma_{\rm max}(\boldsymbol{X})$ produced by the randomized SVD with approximation rank $s=100$. We use $B=1000$ bootstrap iterations and use $1000$ trials to estimate the true standard deviation $\sqrt{\operatorname{Var}(S)}$ of $S$. For this example, the jackknife and bootstrap estimates and true standard deviation were as follows:
$\sqrt{\operatorname{Var}(S))}=8.2\times 10^{-8},\quad\operatorname{Jack}(S)=3.2\times 10^{-7},\quad\mathrm{Boot}(S)=5.0\times 10^{-3}.$
The jackknife overestimates the standard deviation by a factor of 3.9; the bootstrap overestimates the standard deviation by a factor of 60,000.
Existing applications of bootstrap methodology to randomized matrix computations has focused on Monte Carlo-type matrix estimators, a setting in which the bootstrap is more natural. Unfortunately, these Monte Carlo estimators have poor accuracy compared to the randomized SVD. Based on the failure mode presented here, we believe it is unlikely that a generally effective and computationally efficient (i.e., requiring no additional matrix–vector products with $\boldsymbol{A}$) variance estimate based on bootstrap resampling can be developed for the randomized SVD.
Section SM3 Additional numerical experiments
@@
  • Kernel. The QM9 kernel matrix from Section 1.1.
@@
Figure SM2: Leave-one-out error estimator for randomized Nyström. Error and error estimate for randomized SVD ($q=0$) approximation. Includes more test matrices than Fig. 3.
@@
Figure SM3: Matrix jackknife for projectors onto singular subspaces. Error, standard deviation, and jackknife estimate for randomized SVD ($q=0$) approximation $\boldsymbol{X}$ Eq. 19 to the projector $\boldsymbol{\Pi}$ onto the span of the five dominant right singular vectors. Includes more test matrices than Fig. 4.
@@
Section SM4 MATLAB implementations
In this section, we provide MATLAB R2022b implementations of the leave-one-out error estimate and jackknife variance estimate for the Nyström approximation (Section SM4.1). and randomized SVD (Section SM4.2).
SM4.1 Nyström approximation